Session: Feb 1998 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Mantovani at Tue Feb 24 02:44:52 MET 1998) Success
Cloudy during all the observing time.

Torun (K. Borkowski at Tue Feb 24 09:31:23 MET 1998) Success
No known problems

Noto (C.Trigilio at Wed Feb 25 14:36:16 MET 1998) Success

Onsala (Berusad Antonis at Thu Feb 26 13:17:19 MET 1998) Minor failures
Well maybe more than minor failures. The inchword motor for HEAD 2 is broken and the head is not positioned properly. HEAD 1 appears to be OK. However due to the problem in HEAD 2 the pass command timed out and it is possible that during the fists 10 passes HEAD 1 was not always positioned properly. This was corrected by the operator manually from pass 10 onward. Hence the first 3.5 hours may have been lost.

Jodrell (Alastair Gunn at Thu Feb 26 14:56:30 MET 1998) Minor failures
Strong RFI - receiver often saturated. Schedule stopped at 00:09 UT and restarted at scan 11F at 00:18 due to two non-consecutive times in schedule. No other known problems.

Calibration (jean\-francois Desmurs at Mon Mar 2 18:51:47 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the feb98 directory under the name v061c1.antab. Read the comments in the file to get further information.

Effelsberg (dgraham at Thu Mar 5 16:11:37 MET 1998) Success